OpenStudy (anonymous):

Find the inverse function of f(x)=14 + root(3)(x).

OpenStudy (anonymous):

y = 14 + root(3)(x) x = 14 + root(3)(x)

OpenStudy (anonymous):

and then I would cube both sides?

myininaya (myininaya):

did you mean to replace that one x with y?

OpenStudy (anonymous):

Oh, yes. I have to find f^-1(x)

myininaya (myininaya):

right isolate the cuberoot(y) part first by subtracting that other number on both sides then you can talk about cubing both sides

OpenStudy (anonymous):

x-14 = root(3)(y)

OpenStudy (anonymous):

(x-14)^3 = y

myininaya (myininaya):

right

OpenStudy (anonymous):

f^-1(x) = (x-14)^3

myininaya (myininaya):

that is right that is the inverse function of f

OpenStudy (anonymous):

thanks for the help

myininaya (myininaya):

i didn't do anything you did it all by yourself congrats! :)